# Towards Automatic Evaluation for Image Transcreation

### v2

@@ Abstract
Beyond conventional paradigms of translating speech and text, recently, there has been interest in automated transcreation of images to facilitate localization of visual content across different cultures. Attempts to define this as a formal Machine Learning (ML) problem have been impeded by the lack of automatic evaluation mechanisms, with previous work relying solely on human evaluation. In this paper, we seek to close this gap by proposing a suite of automatic evaluation metrics inspired by machine translation (MT) metrics, categorized into: a) Object-based, b) Embedding-based, and c) VLM-based. Drawing on theories from translation studies and real-world transcreation practices, we identify three critical dimensions of image transcreation: cultural relevance, semantic equivalence and visual similarity, and design our metrics to evaluate systems along these axes. Our results show that proprietary VLMs best identify cultural relevance and semantic equivalence, while vision-encoder representations are adept at measuring visual similarity. Meta-evaluation across 7 countries shows our metrics agree strongly with human ratings, with average segment-level correlations ranging from 0.55-0.87. Finally, through a discussion of the merits and demerits of each metric, we offer a robust framework for automated image transcreation evaluation, grounded in both theoretical foundations and practical application. Our code can be found here. [^1] https://github.com/simran-khanuja/automatic-eval-transcreation
@@ 1 Introduction
In this paper, we make progress towards building and evaluating automated metrics for image transcreation. First, given that this is a novel task in ML, we turn to translation studies for a better understanding of its scope and variables (Section 2). We identify key dimensions critical to define the image transcreation process: cultural relevance, semantic equivalence, and visual similarity, and look at real-world applications of transcreation through the lens of these dimensions in Section 3. Grounded in the conceptual and practical insights from previous sections, and drawing inspiration from MT metrics for text, we propose three metrics that account for the varied dimensions of this subjective task – a) Object-based; b) Embedding-based and c) VLM-based metrics, each of which leverages powerful open and closed VLMs and Large Language Models (LLMs). After conducting meta-evaluation along all three dimensions, we report average segment-level (i.e. instance-level) correlations with human ratings of up to 0.87 on image transcreation test-sets spanning 7 countries. Thus, our primary contributions are:
@@ 2 Transcreation in Translation Studies
Parameterizing Transcreation: Broadly, as Venuti (2017) discusses, translation often navigates between the dichotomy of foreignization, which emphasizes preserving the source culture, and domestication, which adapts the content to fit the target culture. Various terms have been used to describe the ends of this spectrum, such as “literal” vs. “free”, “formal” vs. “dynamic”, and “semantic” vs. “communicative”. Yet, no translation decision is ever solely one or the other Venuti (2017) and translators continuously navigate this spectrum, weighing several critical factors.
In transcreation, this dichotomy is even more pronounced, and there is a strong emphasis on making the content feel relevant and engaging in a new cultural context, especially in areas like marketing and entertainment. This often involves adjusting or replacing Culturally Specific Items (CSIs), where cultural references from the source material may not translate well into the target culture Aixelá (1996). The goal is to make the adapted content not only culturally relevant, but also effective in reaching and persuading the audience.
A key factor in guiding the transcreation beyond cultural relevance, is the user/application’s intent behind the process. Skopos theory Reiss et al. (2014) emphasizes that all translation activities are intrinsically purpose-driven. In the context of transcreation, the intent can be as broad as the overall goals of a marketing campaign or as specific as the instructions in a project brief Díaz-Millón and Olvera-Lobo (2023). Whether the aim is to inform, persuade, or entertain, the intent determines how much of the original material is adapted to suit the target audience’s needs Ray and Kelly (2010).
Finally, transcreation requires a high degree of creativity. Creativity is the core component that drives the transcreation process, extending beyond accuracy of faithfulness to the source, to involve a comprehensive reimagining of content Martín (2024). Creativity becomes the mechanism through which intent and cultural relevance are aligned. As Nord (2009) points out, this creative process often demands unique, non-replicable solutions tailored to the specific needs of each project.
@@ 3.2 Comics
Overall, we note that while readers typically prefer transcreations that remain true to the original, adaptations are sometimes necessary to meet social norms, legalities, or cultural differences. Hence, models that pay more attention to semantic equivalence and visual similarity would be preferred, even at the cost of lower cultural relevance for the target audience.
@@ 4 Proposing Automatic Metrics
Drawing from the above lineage, a parallel evolution of metrics for image transcreation is listed below:
@@ 4.1 Object-based metrics
  • Step 2: Obtain Valid Translations: Conditioned on the intent of transcreation, identify the culturally-specific items (CSIs) Aixelá (1996) that would be unfamiliar to the target culture. Subsequently, determine a valid set of replacements for each CSI based on the constraints of this intent.
  • Step 3: Metric Calculation: Calculate the proportion of CSIs that have been correctly replaced. This is determined by checking whether the target CSI is a valid replacement of the source CSI according to Step 2.
@@ 4.2 Embedding-based Metrics
where $\Delta S_{\text{c-r}}$ represents the change in similarity that indicates cultural relevance. $V_{\text{tgt}}$ and $V_{\text{src}}$ are the visual embeddings of the transcreated (target) image and the source image, respectively. $T_{\text{cul}}$ is a textual representation of the cultural reference, such as "This image is from region X.". A higher increase in $\Delta S_{\text{c-r}}$ indicates a greater change towards cultural relevance for the target demographic.
@@ 4.3 VLM-based Metrics
5 Experiments
In this section, we conduct a meta-evaluation of the transcreation metrics proposed above.
@@ 5.1 Evaluation Dataset
Khanuja et al. (2024) create a dataset comprising of commonly occurring concepts across 17 semantic categories (like food, beverages, mammals etc.) that are cross-culturally comprehensive, collected across seven countries. The final dataset includes 585 images, and they apply their proposed pipeline with each country as the target, using images from all other countries as the source. Around 500 images are transcreated for each target country (Table 1). Details about this dataset can be found here. [^4] https://machine-transcreation.github.io/image-transcreation/
5.2 Systems
Khanuja et al. (2024) propose three systems comprising of state-of-the-art LLMs, retrievers and image-editing models. Briefly, the first system, e2e-instruct, is a trivial end-to-end image-editing model prompted to make the image culturally relevant to X, where X is the target country. The second pipeline, cap-edit, first captions the image, then makes the caption ‘culturally relevant’ using a LLM, and finally edits the original image with the culturally-modified caption. The third system, cap-retrieve uses the culturally-modified caption from cap-edit to retrieve a natural image using the modified caption as the query. In their dataset, all outputs are given a rating on a scale of 1-5 by human annotators, across multiple dimensions including visual change, semantic equivalence, cultural relevance, naturalness and offensiveness. They find that the cap-retrieve pipeline performs best, across countries. Questions relevant to our meta-evaluation are in Table 2.
@@ 5.2 Systems
5.3 Correlations with Human Ratings
To evaluate the effectiveness of metrics proposed in Section 4, we calculate segment-level correlations of these metrics with the human ratings from Khanuja et al. (2024) for cultural relevance, semantic equivalence, and visual similarity. Similar to MT that defines segment-level correlation on the order of test instances (sentences), we define it at the level of each image in our test set. We calculate segment-level correlation using the default SciPy (v1.11.4) implementation of Kendall’s Tau (Kendall, 1938), and average across the entire test set for a given country. Segments with non-parsable model outputs, or those scored identically by all systems (usually a small minority) are skipped. We detail how the automated metrics are implemented below:
@@ 5.4 Results
Table 3 presents the average segment-level correlations across seven countries for the three evaluation criteria. For semantic equivalence, proprietary VLMs such as GPT-4o and Gemini-1.5-Pro show the highest performance, likely due to their advanced multimodal reasoning capabilities. We also note that embedding-based metrics outperform open-source VLMs, thus proving to be more cost-effective in budget-lean scenarios. However, as open-source VLMs improve, these dynamics could shift in favor of open-source VLMs. For visual similarity, embedding-based metrics using the SigLIP model excel, achieving a strong correlation of 0.87, significantly surpassing both open-source and proprietary VLMs. This suggests that embedding models are better at capturing fine-grained visual nuances than VLMs that primarily focus on image-text alignment during training. For cultural relevance, the trends mirror those of semantic equivalence, though with lower overall correlations, highlighting the complexity of assessing cultural relevance and the potential for future research in building metrics that capture this dimension.
@@
Table 3: Average segment-level correlation between human ratings and the three automated metrics proposed. ‘cult-rel’ – cultural relevance, ‘sem-eq’ – semantic equivalence, ‘vis-sim’ – visual similarity,
@@ 5.5 Strengths, Limitations, Recommendations
We are inspired by recent research on meta-evaluation in MT that has shown that no single MT metric is effective in capturing all kinds of translation errors and each has its strengths and weaknesses, making a singular overall score from a ‘winning’ metric uninformative (Moghe et al., 2024). We argue that this problem is exacerbated in transcreation, where the open-ended and application-dependent nature of the problem might make an overall score ambiguous. Thus, we now analyze the strengths and limitations of each metric and provide recommendations for the metric models that researchers might use in different scenarios. We also qualitatively analyze the differences across categories in the transcreation dataset that contains 17 categories ranging from concrete (e.g., fruits, birds, beverages), to abstract (e.g. education, religion) ones. Category-wise correlation scores with human ratings can be found here. [^5] https://github.com/simran-khanuja/automatic-eval-transcreation/tree/main/category-correlations
@@ 6 Conclusion
In this paper, we take a first step towards automatically evaluating image transcreation - the task of adapting visual content from a source culture to a target culture. We first review related works in translation studies and identify that cultural relevance for the target audience, semantic equivalence of the transcreated image, and its visual similarity with the source image are three measurable dimensions which potentially cover a wide range of applications. We explore how visual content is being transcreated for real-world applications through this lens, and design evaluation metrics that rank systems across these three axes, drawing inspiration from MT metrics. We find that VLM-based metrics excel at multimodal reasoning (cultural relevance and semantic equivalence) but visual similarities can be better captured using visual-encoder embeddings. For all metrics, capturing cultural relevance of an image for a target demographic is challenging, leaving headroom for progress. We hope that our study helps ML researchers make progress on this important, real-world task.

### v3

@@ Abstract
Beyond conventional paradigms of translating speech and text, recently, there has been interest in automated transcreation of images to facilitate localization of visual content across different cultures. Attempts to define this as a formal Machine Learning (ML) problem have been impeded by the lack of automatic evaluation mechanisms, with previous work relying solely on human evaluation. In this paper, we seek to close this gap by proposing a suite of automatic evaluation metrics inspired by machine translation (MT) metrics, categorized into: a) Object-based, b) Embedding-based, and c) VLM-based. Drawing on theories from translation studies and real-world transcreation practices, we identify three critical dimensions of image transcreation: cultural relevance, semantic equivalence and visual similarity, and design our metrics to evaluate systems along these axes. Our results show that proprietary VLMs best identify cultural relevance and semantic equivalence, while vision-encoder representations are adept at measuring visual similarity. Meta-evaluation across 7 countries shows our metrics agree strongly with human ratings, with average segment-level correlations ranging from 0.55-0.87. Finally, through a discussion of the merits and demerits of each metric, we offer a robust framework for automated image transcreation evaluation, grounded in both theoretical foundations and practical application. Our code can be found here. [^1] https://github.com/simran-khanuja/automatic-eval-img-transcreation
@@ 1 Introduction
In this paper, we make progress towards building and evaluating automated metrics for image transcreation. First, given that this is a novel task in ML, we turn to translation studies for a better understanding of its scope and variables (Section 2). We identify key dimensions critical to define the image transcreation process: cultural relevance, semantic equivalence, and visual similarity, and look at real-world applications of transcreation through the lens of these dimensions in Section 3. Grounded in the conceptual and practical insights from previous sections, and drawing inspiration from MT metrics for text, we propose three metrics that account for the varied dimensions of this subjective task – a) Object-based; b) Embedding-based and c) VLM-based metrics, each of which leverages powerful open and closed VLMs and Large Language Models (LLMs). After conducting meta-evaluation along all three dimensions, we report average segment-level (i.e. instance-level) correlations with human ratings ranging from 0.55-0.87 on image transcreation test-sets spanning 7 countries. Thus, our primary contributions are:
@@ 2 Transcreation in Translation Studies
Parameterizing Transcreation: Broadly, as Venuti (2017) discusses, translation often navigates between the dichotomy of “foreignization”, which emphasizes preserving the source culture, and “domestication”, which adapts the content to fit the target culture. Various terms have been used to describe the ends of this spectrum, such as “literal” vs. “free”, “formal” vs. “dynamic”, and “semantic” vs. “communicative”. Yet, no translation decision is ever solely one or the other Venuti (2017) and translators continuously navigate this spectrum, weighing several critical factors.
In transcreation, this dichotomy is even more pronounced, and there is a strong emphasis on making the content feel relevant and engaging in a new cultural context, especially in areas like marketing and entertainment. This often requires modifying or replacing Culturally Specific Items (CSIs) when source-cultural references do not easily transfer to the target audience Aixelá (1996). The goal is not just to ensure cultural relevance but also to maintain the intended impact of the original content. Thus, every transcreator must strike a balance between preserving semantic equivalence – ensuring the intended meaning remains intact, and achieving cultural relevance by adapting the content for the target audience.
A crucial factor in determining this balance is the intent behind the transcreation, as emphasized by Skopos theory Reiss et al. (2014). The intent can be as broad as the overall goals of a marketing campaign or as specific as the instructions in a project brief Díaz-Millón and Olvera-Lobo (2023). Whether the purpose is to inform, persuade, or entertain, this intent dictates the degree to which the original content should be adapted Ray and Kelly (2010).
Finally, transcreation is inherently a creative process. Beyond fidelity to the source, it demands an imaginative reworking of content to achieve cultural resonance Martín (2024). “Creativity” is the mechanism through which cultural relevance can be achieved. As Nord (2009) points out, this creative process often demands unique, non-replicable solutions tailored to the specific needs of each project. While creativity itself may not be easily quantifiable, evaluating visual similarity between the original and transcreated content provides an indirect means of assessing how faithfully the adaptation retains key visual elements while achieving its intended cultural and semantic transformation.
@@ 3.2 Comics
Overall, we note that while readers typically prefer transcreations that remain true to the original, adaptations are sometimes necessary to meet social norms, legalities, or cultural differences. Hence, models that pay more attention to semantic equivalence and visual similarity would be preferred, even at the cost of lower overall cultural relevance for the target audience.
@@ 4 Proposing Automatic Metrics
Drawing from the above lineage, a parallel evolution of metrics for image transcreation is conceptualised below:
@@ 4.1 Object-based metrics
  • Step 2: Obtain Valid Translations: Conditioned on the intent of transcreation, identify the culturally-specific items (CSIs) Aixelá (1996) in the source image, that would be unfamiliar to the target culture. Subsequently, determine a valid set of replacements for each CSI based on the constraints of this intent.
  • Step 3: Metric Calculation: Calculate the proportion of CSIs that have been correctly replaced in the model output. This is determined by checking whether objects identified in the target image are valid replacements of the source CSIs according to Step 2.
@@ 4.2 Embedding-based Metrics
where $\Delta S_{\text{c-r}}$ represents the change in similarity that indicates cultural relevance. A higher increase in $\Delta S_{\text{c-r}}$ indicates a greater change towards cultural relevance for the target demographic.
@@ 4.3 VLM-based Metrics
5 Experimental Setup
In this section, we perform a meta-evaluation of the transcreation metrics introduced above. Specifically, we assess the outputs of transcreation systems proposed in Khanuja et al. (2024) using the metrics defined in Section 4 and analyze the correlation between our scores and human ratings for the same system outputs. [^4] Available at: https://github.com/simran-khanuja/image-transcreation/tree/main/human_evaluation
@@ 5.1 Evaluation Dataset
Khanuja et al. (2024) construct a dataset featuring commonly occurring concepts across 17 semantic categories (such as food, beverages, and mammals) that are culturally comprehensive, with data collected from seven different countries. The final source dataset consists of 585 images. Each of the seven countries serves as a target, with source images drawn from the remaining six countries, resulting in approximately 500 images being transcreated for each target country (Table 1). Further details on this dataset can be found here. [^5] https://machine-transcreation.github.io/image-transcreation/ The task involves converting an image into another within the same semantic category while ensuring greater cultural relevance for a specific target country, X. For example, a food item in the source image may be replaced with a different food item that is more commonly recognized or significant in X.
5.2 System Descriptions
Khanuja et al. (2024) propose three systems comprising of state-of-the-art LLMs, retrievers and image-editing models for the image transcreation task. Briefly, the first system, e2e-instruct, is a trivial end-to-end image-editing model prompted to make the image culturally relevant to X, where X is the target country. The second pipeline, cap-edit, first captions the image, then makes the caption ‘culturally relevant’ using a LLM, and finally edits the original image with the culturally-modified caption. The third system, cap-retrieve uses the culturally-modified caption from cap-edit to retrieve a natural image using the modified caption as the query. In their dataset, all outputs are given a rating on a scale of 1-5 by human annotators, across multiple dimensions including visual change, semantic equivalence, cultural relevance, naturalness and offensiveness. They find that the cap-retrieve pipeline performs best, across countries. Questions relevant to our meta-evaluation are in Table 2.
@@ 5.2 System Descriptions
5.3 Metric Implementation
We calculate segment-level correlations of proposed metric scores with human ratings. Similar to MT that defines segment-level correlation on the order of test instances (sentences), we define it at the level of each image in our test set. We use the default SciPy (v1.11.4) implementation of Kendall’s Tau (Kendall, 1938), and average across the entire test set for a given country. Segments with non-parsable model outputs, or those scored identically by all systems (usually a small minority) are skipped. Our metrics are implemented as follows:
@@ 5.4 Results
Table 3 presents the average segment-level correlations across seven countries for the three evaluation criteria.
For semantic equivalence, proprietary VLMs such as GPT-4o and Gemini-1.5-Pro show the highest performance, likely due to their advanced multimodal reasoning capabilities. We also note that embedding-based metrics outperform open-source VLMs, thus proving to be more cost-effective in budget-lean scenarios. However, as open-source VLMs improve, these dynamics could shift in favor of open-source VLMs.
For visual similarity, embedding-based metrics using the SigLIP model excel, achieving a strong correlation of 0.87, significantly surpassing both open-source and proprietary VLMs. This suggests that embedding models are better at capturing fine-grained visual nuances than VLMs that primarily focus on image-text alignment during training.
For cultural relevance, the trends mirror those of semantic equivalence, though with lower overall correlations, highlighting the complexity of assessing cultural relevance and the potential for future research in building metrics that capture this dimension.
@@
Table 3: Average segment-level correlation between human ratings and the three automated metrics proposed. ‘cult-rel’ – cultural relevance, ‘sem-eq’ – semantic equivalence, ‘vis-sim’ – visual similarity.
@@ 5.5 Strengths, Limitations, Recommendations
We are inspired by recent research on meta-evaluation in MT that has shown that no single MT metric is effective in capturing all kinds of translation errors and each has its strengths and weaknesses, making a singular overall score from a ‘winning’ metric uninformative (Moghe et al., 2024). We argue that this problem is exacerbated in transcreation, where the open-ended and application-dependent nature of the problem might make an overall score ambiguous. Thus, we now analyze the strengths and limitations of each metric and provide recommendations for the metric models that researchers might use in different scenarios. Section 3 is an attempt at elucidating the axes of transcreation which can potentially be more or less relevant for a given application. We also qualitatively analyze the differences across categories in the transcreation dataset that contains 17 categories ranging from concrete (e.g., fruits, birds, beverages), to abstract (e.g. education, religion) ones. Category-wise correlation scores with human ratings can be found here. [^6] https://github.com/simran-khanuja/automatic-eval-img-transcreation/tree/main/category-correlations
@@ 6 Conclusion
In this paper, we take a first step towards automatically evaluating image transcreation - the task of adapting visual content from a source culture to a target culture. We first review related works in translation studies and identify that cultural relevance for the target audience, semantic equivalence of the transcreated image, and its visual similarity with the source image are three measurable dimensions which potentially cover a wide range of applications. We explore how visual content is being transcreated for real-world applications through this lens, and design evaluation metrics that rank systems across these three axes, drawing inspiration from MT metrics. We find that VLM-based metrics excel at measuring cultural relevance and semantic equivalence, but visual similarities can be better captured using visual-encoder embeddings. For all metrics, capturing cultural relevance of an image for a target demographic is challenging, leaving headroom for progress. We hope that our study helps ML researchers make progress on this important, real-world task.
@@ 9 Ethical Statement
10 Contributions
We list author contributions loosely following the CRediT author statement. [^7] https://www.elsevier.com/researcher/author/policies-and-guidelines/credit-author-statement
  • SK: Conceptualization (co-lead), Methodology (lead), Software (co-lead), Resources (lead), Writing (co-lead), Investigation (co-lead), Visualization (co-lead)
  • VI: Conceptualization (co-lead), Software (co-lead), Writing (co-lead), Investigation (co-lead), Visualization (co-lead)
  • CH: Writing (supporting), Investigation (supporting), Resources (supporting)
  • GN: Supervision (lead), Project Administration (lead), Funding Acquisition (lead), Writing (supporting)